| Operation Summary for February 20, 2012 - February 22, 2012 |
|
Wednesday February 22 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operations halted the MI ramp at 9:23 AM to allow Power Supply personnel to inspect the MI-30 quadrupole magnet power supply; everything was normal. MI resumed sending beam at 9:52 AM. A Booster person began a study at 10:02 AM. At 10:07 AM, a Booster bulk power supply (B:CEP400) tripped off for the first of two times in 2.5 hours; it reset. Operators searched and secured the Meson MTest enclosures and began sending beam to the Excellence in Detector Instrumentation Technology (EDIT) School at 11:31 AM. Water Group personnel put the MI LCW system into recirculation mode to reconfirm that the MI LCW magnet system leak goes away. At 2:37 PM, Water Group personnel verified that the water loss from the MI magnet LCW system had stopped after it was taken off of CUB. At 3:32 PM, a Linac RF station (LRF1) tripped off; it reset. During the evening shift, at 4:51 PM, Operators reported problems with a Switchyard (SY) CAMAC crate; Operators investigated and reported some odd things, but since beam was making it through to the EDIT School decided not to do anything until later; later the problem went away. The Booster study ended at 4:53 PM. Operations disabled beam to the MTest EDIT School at 5:53 PM. At 6:49 PM, a Booster RF station (BRF19) tripped off for the first of four times in about one hour; it reset. Operations reported making some significant phase changes to MI at 9:46 PM. BRF14 and BRF15 tripped off due to sparking; they both reset. Thursday February 23 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, BRF11 tripped off at 1:08 AM; it reset. At 3:22 AM, BRF15 and 16 tripped off; they both reset. BRF2 tripped off at 5:19 AM; it reset. At 5:32 AM, BRF12 tripped off; it reset. At 7:28 AM, Neutron Therapy Facility (NTF) personnel prepared to treat three patients during the morning. Operators reported at 7:38 AM that they had completed a 750 KeV line tuning and Booster injection and capture tuning during the shift. During the day shift, Operators reported at 8:18 AM that they had made MI phase adjustments. B:CEP400 tripped off at 9:53 AM and also brought down a quadrupole magnet power supply (Q803); Power Supply personnel investigated and brought them both back online. BRF3 tripped off at 10:46 AM; it reset. Water Group personnel reported at 11:04 AM that they were putting the MI LCW system in its normal circulation mode with CUB to continue their search for the MI LCW leak. Operators loaded protons into the Recycler at 1:21 PM. Operators completed a search and secure of the Neutrino Muon Lab (NML) at 2:05 PM. Starting around 2:43 PM, SY personnel began splitting some of the Meson beam to the MCenter dump. During the evening shift, around 7:15 PM, the FTBF Meson experiment T-979 began taking beam. Friday February 24 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and the FTBF experiment T-979. During the midnight shift, Operations disabled beam to T-979 at 12:07 AM. At 1:52 AM, the MI South Anode Power Supply tripped off on an interlock indication; it reset. BRF3 tripped off at 2:13 AM; it reset. At 2:16 AM, BFR16 tripped off; it reset. Operators reported at 5:48 AM that they had adjusted the MI tunes. At 6:45 AM, three Linac quadrupole magnet power supplies (QPS104, 105, and 108) tripped off; they all reset. At 7 AM, NTF personnel prepared to treat three patients during the morning. At 7:07 AM, the H- Source tripped off; it reset. At 7:19 AM, a Booster vacuum pump (B:IPRF22) tripped off and brought down BRF11 and 12; everything recovered and reset. Operators reported that they had conducted Booster injection tuning throughout the shift. Operations Coordinator (OpsCo) Plans for Today and the Weekend The OpsCo wants the complex to operate with as few beam disruptions as possible for the EDIT School and the other experiments. The EDIT School ends today (2/24/12) at 6 PM. |
|
More Information Comments and Suggestions |
| last modified 02/24/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |